# CHIPS-FF: Evaluating Universal Machine Learning Force Fields for Material Properties

### v1

@@ Abstract
In this work, we introduce CHIPS-FF (Computational High-Performance Infrastructure for Predictive Simulation-based Force Fields), a universal, open-source benchmarking platform for machine learning force fields (MLFFs). This platform provides robust evaluation beyond conventional metrics such as energy, focusing on complex properties including elastic constants, phonon spectra, defect formation energies, surface energies, and interfacial and amorphous phase properties. Utilizing 13 graph-based MLFF models including ALIGNN-FF, CHGNet, MatGL, MACE, SevenNet, ORB and OMat24, the CHIPS-FF workflow integrates the Atomic Simulation Environment (ASE) with JARVIS-Tools to facilitate automated high-throughput simulations. Our framework is tested on a set of 104 materials, including metals, semiconductors and insulators representative of those used in semiconductor components, with each MLFF evaluated for convergence, accuracy, and computational cost. Additionally, we evaluate the force-prediction accuracy of these models for close to 2 million atomic structures. By offering a streamlined, flexible benchmarking infrastructure, CHIPS-FF aims to guide the development and deployment of MLFFs for real-world semiconductor applications, bridging the gap between quantum mechanical simulations and large-scale device modeling.
@@
MACE68, 69, 54, an equivariant message passing neural network potential, uses higher order messages (higher than two body) combined with atomic cluster expansion (ACE) 70, 71, which is a method for deriving an efficient body-ordered symmetric polynomial basis to represent functions of atomic neighborhoods. A pre-trained model for the entire periodic table (trained on MPTrj) for MACE was released in late 2023 54 (referred to as MACE-MP-0). SevenNet (Scalable EquiVariance Enabled Neural Network) 72, another equivariant message passing uMLFF, is based on the NequIP 73 architecture for building E(3)-equivariant force fields. SevenNet trained on the MPtrj dataset and was released in 2024. In addition to these open source models, there have been proprietary uMLFF models developed including MatterSim (Microsoft) 74 and GNoME (Google DeepMind) 75. A version of MatterSim was made open source to the community in December 2024 76. The Orb 77 pretrained uMLFF, released in fall 2024, was developed by the startup company Orbital Materials, where superior performance for force and energy predictions were demonstrated at significantly reduced computational cost 55, 78, 79. Orb utilizes an attention augmented Graph Network-based Simulator (GNS) 80, which is a type of Message Passing Neural Network (MPNN) 81. In contrast to equivariant models such as MACE and SevenNet, Orb does not utilize equivariant message passing. The open source Orb model (orb-v2) is trained on MPtrj and Alexandria (over 3 million materials and 32 million relaxation steps) 55. In late 2024, Meta FAIR Chemistry released the Open Materials 2024 (OMat24) 82 models and dataset, which contains over 110 million DFT calculations and pretrained uMLFF models based on the EquiformerV2 83 architecture, which is a highly accurate equivariant transformer model. OMat24 offers various sized models trained on OMat, MPTrj and a subset of Alexandria.
@@
Recently, interest in the field of uMLFFs has grown immensely (from academia, government and industry), as seen by the several models above and various use cases of new model architectures. Given the current state of the field, benchmarking the performance of uMLLFs is imperative. Large-scale efforts to benchmark the performance of various pretrained uMLFF architectures have emerged such as Matbench Discovery 84, 79 and the JARVIS-Leaderboard 85, 86, which calculate error metrics and allow for new models to be uploaded as time progresses. There have beenfocused benchmarking efforts for more involved properties beyond energy. Yu et al. 87 performed a systematic assessment of uMLFFs (M3GNET, ALIGNN-FF, CHGNet and MACE), calculating properties such as optimized structure, formation energy, bulk modulus, and phonon band structure, comparing directly to DFT to calculate error metrics. Benchmarking of more advanced properties such as surface energies, defect formation energies, solid-solution energetics, and ion migration barriers have recently been performed with M3GNET, CHGNet and MACE 46, 47. In addition, a high-level assessment of uMLFF computed thermal conductivity has been conducted with M3GNET, CHGNet, MACE, SevenNet and ORB by Pot́a et al. 51 More recently, Zhu et al. utilized M3GNet, CHGNet, MACE, SevenNet, and ORB to accelerate the prediction of CALPHAD-based phase diagrams for complex alloys 88. These recent efforts have highlighted the successes and limitations of various uMLFF architectures, and have emphasized the need for a comprehensive publicly available benchmarking platform. Such a platform allows the community to make informed decisions on regarding uMLFFs, outweighing factors such as computational cost versus accuracy. In addition, researchers may find systematic trends in the uMLFF results which can help others to understand the applicability of each uMLFF type. Well-understood systematic trends in DFT functionals (i.e., the underbinding and lattice constant errors of GGA) have been exploited to identify new exfoliable van der Waals (vdW) materials 89. In contrast to the Jacob’s Ladder metaphor of DFT exchange-correlation functionals 90, where the “rungs” of the ladder refer to the increasing complexity (and computational cost) and accuracy of the functionals, there (as of yet) exists no systematic assessment or generalizable statements for uMLFFs. As uMLFFs begin to replace DFT calculations, such analysis is needed to help the community better develop new force fields or fine-tune existing force fields for downstream applications.
@@
In this work, we present CHIPS-FF (Computational High-performance Infrastructure for Predictive Simulation-based Force Fields): a generalized user-friendly open-source benchmarking package to test various uMLFFs for a number of properties beyond energy. We have implemented a streamlined workflow to use uMLFFs for structural relaxation, bulk modulus, elastic properties, point defect formation energy, surface energy, interfacial properties (work of adhesion), molecular dynamics and creation of amorphous structures, all with automated error metrics from JARVIS-DFT. In contrast to large scale benchmarking efforts such as Matbench Discovery 84, 79, which requires users to submit contributions covering enormous test sets (on the order of 200,000), CHIPS-FF allows for robust benchmarking on smaller datasets and more complex properties. Although our framework is agnostic to the type of material, we chose a set of 104 materials that are most relevant for semiconductor devices as a case study. Additionally, we evaluate force-prediction accuracy of these models for close to 2 million atomic structures from Materials project and JARVIS-DFT databases. We intend this codebase and benchmarking dataset to benefit the MLFF community and aid in the development of future models.
The CHIPS-FF workflow mainly utilizes the Atomic Simulation Environment (ASE) 91 and JARVIS-Tools 38, 39 to perform atomistic simulations in Python. In terms of pretrained uMLFF models, we utilized ALIGNN-FF: 2.12.2024 (trained on 307k entries from JARVIS-DFT), CHGNet: 0.3.8, , MatGL: 1.1.2 (using both the M3GNet-MP-2021.2.8-PES and M3GNet-MP-2021.2.8-DIRECT-PES 92 models), MACE-MP-0: 0.3.6, SevenNet-0: 0.9.2 (using the 11July2024 model), and orb-models: 0.4.0 (orb-v2). In addition, we benchmarked a MACE force field trained on the newly released Alexandria dataset presented in Ref. 45 (this model was trained on 400,000 calculations with the PBEsol 93 functional). For the OMat24 models from Meta, we tested the three EquiformerV2 (eqV2) models trained on OMat (Small: 31M (million), medium: 86M, and large: 153M in terms of model parameters) and two eqV2 models trained on OMat+MPTrj+subset of Alexandria (Small: 31M and medium: 86M) 56, 82. These models are named as: eqV2_31M_omat, eqV2_86M_omat, eqV2_153M_omat, eqV2_31M_omat_mp_salex, and eqV2_86M_omat_mp_salex 56, 82.
@@
where $E_{\text{surface}}$ is the energy of the surface structure, $E_{\text{bulk}}$ is the energy of the bulk structure, $N$ is the number of bulk unit cells in the surface, and $A$ is the cross sectional area of the surface. In order to study amorphous materials, we have implemented finite-temperature molecular dynamics into our workflow to perform melt/quench simulations using Berendsen NVT (Number of particles, Volume, Temperature ensemble) dynamics within ASE. For our case study of amorphous Si, we used a timestep of 1 femtosecond and ran the simulation at 3500 K for 10 ps (melt) and 300 K for 20 ps (quench). For benchmarking calculations to compare uMLFF results to, we performed similar ab initio MD (AIMD) calculations with the Vienna Ab initio Simulation Package (VASP) code, using projector augmented wave (PAW) pseudopotentials 102, 103 and the vdW-DF-optB88 functional. In order to perform calculations for material interfaces, we utilized InterMat 21, a newly developed package for the generation and calculation of interface structures (substrate plus film). The initial interface is obtained by creating a superlattice or alternating slab junction (ASJ) structure (without vacuum padding) 104, 105 and then using the Zur algorithm 106 to obtain the best candidate interface. From here, we computed the relative alignment between the film and substrate in the in-plane (xy) direction by performing a grid search with a 0.05 fractional spacing interval with each respective uMLFF model. In addition to determining the optimal in-plane orientation between the substrate and film, this gives us a qualitative estimate of how smooth the potential energy surface is at the interface for each pretrained uMLFF. From this calculation, we also determined the work of adhesion ($W_{\text{ad}}$) at the interface with:
@@
A majority of the reference data used to compare our uMLFF results to were obtained from the JARVIS-DFT database, which contains relaxed structural data (primitive lattice vectors, volume), elastic properties (bulk modulus, elastic tensor) and phonon band structure 38, 39. Within JARVIS-DFT, there exists a database of single vacancy calculations 16 and surfaces 21 that were both used for benchmarking these uMLFFs. All of the DFT data in JARVIS was computed with the vdW-DF-optB88 107 functional, which is a modified Generalized Gradient Aprroximation (GGA) functional designed to explicitly include van der Waals (vdW) interactions. This is slightly different than other materials repositories such as the Materials Project (MPtrj), which mostly contains DFT calculations performed with the Perdew Burke Ernzerhof (PBE) 108 GGA functional. Despite mostly comparing to JARVIS-DFT results (vdW-DF-optB88), the CHIPS-FF workflow has a flexible framework to use any DFT dataset as a ground truth (i.e., the Materials Project, Alexandria, a users own DFT dataset, etc.). In addition, the platform is flexible enough to compare uMLFF results directly to available experimental values.
@@
uMLFF Type Bulk Surface Vacancy alignn_ff 6.7 40.2 39.6 chgnet 0.0 1.2 6.3 eqV2_153M_omat 0.0 15.9 29.2 eqV2_31M_omat 0.0 17.1 20.8 eqV2_31M_omat_mp_salex 0.0 6.1 0.0 eqV2_86M_omat 1.0 22.0 25.0 eqV2_86M_omat_mp_salex 0.0 6.1 2.1 mace 0.0 7.3 27.1 mace-alexandria 2.9 14.6 33.3 matgl 0.0 4.9 14.6 matgl-direct 3.9 6.1 12.5 orb-v2 1.0 3.6 0.0 sevennet 0.0 2.4 41.7
@@
Table 2: The mean absolute error (MAE) and total computational time for a variety of material properties calculated with each uMLFF type. The MAE is calculated with respect to JARVIS-DFT data. Properties include (in order) lattice constants a, b, c, volume, bulk modulus, C11 and C44 components of the elastic tensor, and phonon band structure. The computational time is measured per CPU, with the mace-alexandria model which was run on GPU (indicated by ). An interactive and up-to-date version of the table is available in the JARVIS-Leaderboard.
@@
uMLFF Type err_a err_b err_c err_vol err_kv err_c11 err_c44 err_phonon Time Å Å Å Å3 GPa GPa GPa cm-1 seconds alignn_ff 0.087 0.102 0.137 9.95 118 197 74 132 7.7 x 104 chgnet 0.046 0.049 0.109 3.58 89 55 50 60 3.7 x 104 eqV2_153M_omat 0.036 0.041 0.111 2.87 111 78 35 35 1.7 x 106 eqV2_31M_omat 0.029 0.032 0.099 3.05 113 70 38 35 4.0 x 105 eqV2_31M_omat_mp_salex 0.028 0.030 0.096 3.17 112 49 42 34 3.4 x 105 eqV2_86M_omat 0.035 0.041 0.096 3.44 116 81 36 36 1.3 x 106 eqV2_86M_omat_mp_salex 0.026 0.030 0.100 3.14 113 46 37 34 8.5 x 105 mace 0.035 0.038 0.084 3.00 97 43 42 46 8.4 x 104 mace-alexandria 0.081 0.086 0.206 6.05 68 64 51 65 2.1 x 105* matgl 0.052 0.057 0.128 3.50 77 67 46 67 1.8 x 104 matgl-direct 0.044 0.046 0.106 2.88 77 80 50 75 2.7 x 104 orb-v2 0.023 0.025 0.159 3.38 118 86 74 36 3.5 x 104 sevennet 0.035 0.038 0.100 3.43 107 46 40 42 6.2 x 104
@@
Figure 3: Parity plots for a) vacancy formation energy (in eV) and b) surface energy (in J/m2) for each uMLFF type with respect to JARVIS-DFT data (MAE is given in the inset). c) Depicts the parity plot for work of adhesion (Wad in J/m2) for several material interfaces and MAE with respect to previous calculated and experimental data (taken from Ref. 109). To the right of the parity plot, we focus on the best performing Wad prediction for the Al(111)Al2O3(001) interface, showing the xy grid search and optimal in-plane interface configurations with each uMLFF.
@@
Figure 4: The calculated radial distribution function ($g(r)$) as a function of pair separation distance computed with several uMLFF models for amorphous Si (by performing melt/quench simulations). Ab initio (AIMD) results are given as a ground truth benchmark (black dotted line) and the MAE with respect to AIMD is reported in the inset. The atomic structure for amorphous Si is given for AIMD and the best performing (eqV2_31M_omat) and worst performing (ORB) uMLFF models.
@@
Fig. 1 depicts a full schematic of the CHIPS-FF workflow. In this workflow, initial structures are taken from the JARVIS-DFT database. JARVIS-Tools is used to pull structures from the JARVIS-DFT database and generate supercell surface and defect structures prior to atomistic simulations. Various uMLFF calculators are accessed through ASE. So far, we have implemented MatGL, CHGNet, ALIGNN-FF, MACE, SevenNet, ORB, and eqV2 (OMat24). As new pretrained force fields are developed over time, they will be added to the workflow. In addition, CHIPS-FF offers the flexibility for the user to add their own force field or property to the pipeline. Once each uMLFF is loaded and run through the ASE calculator, a variety of properties can be computed including: optimized structure, bulk modulus (equation of state fitting), elastic tensor (elastic package), phonons (phonopy), thermal expansion (phonopy), thermal conductivity (phono3py), molecular dynamics, vacancy formation energy, surface energy, and interface properties (InterMat). As these calculations are carried out, detailed logging information is saved, including the computational timing for each stage of the calculation and whether or not each stage converged. Once these material properties are computed with each uMLFF, they are cross checked with entries in the JARVIS-DFT database and the mean absolute error (MAE) for each property is computed. For spectral quantities such as the phonon band structure, we computed the MAE for each band along the q-point path using:
@@
In addition to summing the computational timing for the entire workflow, we also examined the scaling behavior of each uMLFF (on CPU). These results for supercells of Cu (up to 10,000 atoms) are depicted in Fig. 2a). Unsurprisingly, we see that OMat model scales much more drastically with system size as compared to the other uMLFF models. In addition, we find that the invariant models have more favorable scaling than the equivariant models such as MACE and SevenNet, with SevenNet having slightly better scaling than MACE. The scalability of the ORB model is a huge advantage due to its accuracy of various material property predictions. Similar scalability studies were conducted in Ref. 55 The quantities of surface energy and vacancy formation energy are difficult to predict with machine learning methods and provide a rigorous benchmarking test for uMLFF architectures. Out of the 104 materials, there exists 82 entries for surface DFT calculations and 48 entries for defect DFT calculations within JARVIS-DFT. We used these calculations to compare our uMLFF results. As previously seen in Table 1, convergence for surface and defect supercells is a much more difficult task. For this reason, we narrowed down our set of surfaces and defect calculations to a common set of structures where all of the uMLFFs were able to converge the calculation (excluding ALIGNN-FF due to its very low rate of convergence). We found that 50 surfaces and 20 defect supercells converged for all of the considered uMLFFs. Fig. 3a) and Fig. 3b) depict the parity plots and corresponding MAE for vacancy formation energy and surface energy of this common set of materials. The very low error for surface energy (0.15 J/m2) and vacancy formation energy (0.67 eV) highlight some of the successes of recent state-of-the-art models such as OMat24 and ORB. We also observe generally good results for MACE and SevenNet when predicting vacancy formation energy and surface energies. The low cost of ORB, high rate of convergence and high accuracy make it a viable tool to relax larger surfaces and defect supercells.

### v3

@@ Abstract
In this work, we introduce CHIPS-FF (Computational High-Performance Infrastructure for Predictive Simulation-based Force Fields), a universal, open-source benchmarking platform for machine learning force fields (MLFFs). This platform provides robust evaluation beyond conventional metrics such as energy, focusing on complex properties including elastic constants, phonon spectra, defect formation energies, surface energies, and interfacial and amorphous phase properties. Utilizing 16 graph-based MLFF models including ALIGNN-FF, CHGNet, MatGL, MACE, SevenNet, ORB, MatterSim and OMat24, the CHIPS-FF workflow integrates the Atomic Simulation Environment (ASE) with JARVIS-Tools to facilitate automated high-throughput simulations. Our framework is tested on a set of 104 materials, including metals, semiconductors and insulators representative of those used in semiconductor components, with each MLFF evaluated for convergence, accuracy, and computational cost. Additionally, we evaluate the force-prediction accuracy of these models for close to 2 million atomic structures. By offering a streamlined, flexible benchmarking infrastructure, CHIPS-FF aims to guide the development and deployment of MLFFs for real-world semiconductor applications, bridging the gap between quantum mechanical simulations and large-scale device modeling.
@@
MACE68, 69, 54, an equivariant message passing neural network potential, uses higher order messages (higher than two body) combined with atomic cluster expansion (ACE) 70, 71, which is a method for deriving an efficient body-ordered symmetric polynomial basis to represent functions of atomic neighborhoods. A pre-trained model for the entire periodic table (trained on MPTrj) for MACE was released in late 2023 54 (referred to as MACE-MP-0). In late 2024, a new pretrained version of MACE (known as MACE-MPA-0) was released, which is trained on a combination of MPTrj and Alexandria 72. SevenNet (Scalable EquiVariance Enabled Neural Network) 73, another equivariant message passing uMLFF, is based on the NequIP 74 architecture for building E(3)-equivariant force fields. SevenNet trained on the MPtrj dataset and was released in 2024. In addition to these open source models, there have been proprietary uMLFF models developed including MatterSim (Microsoft) 75 and GNoME (Google DeepMind) 76. A version of MatterSim (MatterSim-v1), which has a similar but slightly modified architecture to M3GNet and is trained on over 17 million structures sampled from Materials Project, Alexandria, and Microsoft’s own dataset (consisting of additional structures and MD trajectories at ambient and extreme conditions), was made open source to the community in December 2024 77. The Orb 78 pretrained uMLFF, released in fall 2024, was developed by the startup company Orbital Materials, where superior performance for force and energy predictions were demonstrated at significantly reduced computational cost 55, 79, 80. Orb utilizes an attention augmented Graph Network-based Simulator (GNS) 81, which is a type of Message Passing Neural Network (MPNN) 82. In contrast to equivariant models such as MACE and SevenNet, Orb does not utilize equivariant message passing. The open source Orb model (orb-v2) is trained on MPtrj and Alexandria (over 3 million materials and 32 million relaxation steps) 55. In late 2024, Meta FAIR Chemistry released the Open Materials 2024 (OMat24) 83 models and dataset, which contains over 110 million DFT calculations and pretrained uMLFF models based on the EquiformerV2 84 architecture, which is a highly accurate equivariant transformer model. OMat24 offers various sized models trained on OMat, MPTrj and Alexandria.
@@
Recently, interest in the field of uMLFFs has grown immensely (from academia, government and industry), as seen by the several models above and various use cases of new model architectures. Given the current state of the field, benchmarking the performance of uMLLFs is imperative. Large-scale efforts to benchmark the performance of various pretrained uMLFF architectures have emerged such as Matbench Discovery 85, 80 and the JARVIS-Leaderboard 86, 87, which calculate error metrics and allow for new models to be uploaded as time progresses. There have been focused benchmarking efforts for more involved properties beyond energy. Yu et al. 88 performed a systematic assessment of uMLFFs (M3GNet, ALIGNN-FF, CHGNet and MACE), calculating properties such as optimized structure, formation energy, bulk modulus, and phonon band structure, comparing directly to DFT to calculate error metrics. Benchmarking of more advanced properties such as surface energies, defect formation energies, solid-solution energetics, and ion migration barriers have recently been performed with M3GNet, CHGNet and MACE 46, 47. In addition, a high-level assessment of uMLFF computed thermal conductivity has been conducted with M3GNet, CHGNet, MACE, SevenNet and ORB by Pot́a et al. 51 More recently, Zhu et al. utilized M3GNet, CHGNet, MACE, SevenNet, and ORB to accelerate the prediction of CALPHAD-based phase diagrams for complex alloys 89 and Lowe et al. performed in-depth benchmarking of phonons for over 10,000 materials using all of the recent state-of-the-art uMLFFs 90. These recent efforts have highlighted the successes and limitations of various uMLFF architectures, and have emphasized the need for a comprehensive publicly available benchmarking platform. Such a platform allows the community to make informed decisions on regarding uMLFFs, outweighing factors such as computational cost versus accuracy. In addition, researchers may find systematic trends in the uMLFF results which can help others to understand the applicability of each uMLFF type. Well-understood systematic trends in DFT functionals (i.e., the underbinding and lattice constant errors of GGA) have been exploited to identify new exfoliable van der Waals (vdW) materials 91. In contrast to the Jacob’s Ladder metaphor of DFT exchange-correlation functionals 92, where the “rungs” of the ladder refer to the increasing complexity (and computational cost) and accuracy of the functionals, there (as of yet) exists no systematic assessment or generalizable statements for uMLFFs. As uMLFFs begin to replace DFT calculations, such analysis is needed to help the community better develop new force fields or fine-tune existing force fields for downstream applications.
@@
In this work, we present CHIPS-FF (Computational High-performance Infrastructure for Predictive Simulation-based Force Fields): a generalized user-friendly open-source benchmarking package to test various uMLFFs for a number of properties beyond energy. We have implemented a streamlined workflow to use uMLFFs for structural relaxation, bulk modulus, elastic properties, point defect formation energy, surface energy, interfacial properties (work of adhesion), molecular dynamics and creation of amorphous structures, all with automated error metrics from JARVIS-DFT. In contrast to large scale benchmarking efforts such as Matbench Discovery 85, 80, which requires users to submit contributions covering enormous test sets (on the order of 200,000), CHIPS-FF allows for robust benchmarking on smaller datasets and more complex properties. Although our framework is agnostic to the type of material, we chose a set of 104 materials that are most relevant for semiconductor devices as a case study. Additionally, we evaluate force-prediction accuracy of these models for close to 2 million atomic structures from Materials Project and JARVIS-DFT databases. We intend this codebase and benchmarking dataset to benefit the MLFF community and aid in the development of future models.
[…]
@@
where $E_{\text{surface}}$ is the energy of the surface structure, $E_{\text{bulk}}$ is the energy of the bulk structure, $N$ is the number of bulk unit cells in the surface, and $A$ is the cross sectional area of the surface. In order to study amorphous materials, we have implemented finite-temperature molecular dynamics into our workflow to perform melt/quench simulations using Berendsen NVT (Number of particles, Volume, Temperature ensemble) dynamics within ASE. For our case study of amorphous Si, we used a timestep of 1 femtosecond and ran the simulation at 3500 K for 10 ps (melt) and 300 K for 20 ps (quench). For benchmarking calculations to compare uMLFF results to, we performed similar ab initio MD (AIMD) calculations with the Vienna Ab initio Simulation Package (VASP) code, using projector augmented wave (PAW) pseudopotentials 106, 107 and the vdW-DF-optB88 functional. In order to perform calculations for material interfaces, we utilized InterMat 21, a newly developed package for the generation and calculation of interface structures (substrate plus film). The initial interface is obtained by creating a superlattice or alternating slab junction (ASJ) structure (without vacuum padding) 108, 109 and then using the Zur algorithm 110 to obtain the best candidate interface. From here, we computed the relative alignment between the film and substrate in the in-plane (xy) direction by performing a grid search with a 0.05 fractional spacing interval with each respective uMLFF model. In addition to determining the optimal in-plane orientation between the substrate and film, this gives us a qualitative estimate of how smooth the potential energy surface is at the interface for each pretrained uMLFF. From this calculation, we also determined the work of adhesion ($W_{\text{ad}}$) at the interface with:
@@
A majority of the reference data used to compare our uMLFF results to were obtained from the JARVIS-DFT database, which contains relaxed structural data (primitive lattice vectors, volume), elastic properties (bulk modulus, elastic tensor) and phonon band structure 38, 39. Within JARVIS-DFT, there exists a database of single vacancy calculations 16 and surfaces 21 that were both used for benchmarking these uMLFFs. All of the DFT data in JARVIS was computed with the vdW-DF-optB88 111 functional, which is a modified Generalized Gradient Aprroximation (GGA) functional designed to explicitly include van der Waals (vdW) interactions. This is slightly different than other materials repositories such as the Materials Project (MPtrj), which mostly contains DFT calculations performed with the Perdew Burke Ernzerhof (PBE) 112 GGA functional. Despite mostly comparing to JARVIS-DFT results (vdW-DF-optB88), the CHIPS-FF workflow has a flexible framework to use any DFT dataset as a ground truth (i.e., the Materials Project, Alexandria, a users own DFT dataset, etc.). In addition, the platform is flexible enough to compare uMLFF results directly to available experimental values.
@@
uMLFF Type Bulk Surface Vacancy alignn_ff 6.7 37.5 39.6 chgnet 0.0 1.1 6.2 eqV2_153M_omat 0.0 14.8 31.2 eqV2_31M_omat 0.0 17.0 25.0 eqV2_31M_omat_mp_salex 0.0 5.7 0.0 eqV2_86M_omat 1.9 18.2 27.1 eqV2_86M_omat_mp_salex 0.0 5.7 2.1 mace 0.0 6.8 25.0 mace-alexandria 2.9 15.9 33.3 mace-mpa 0.0 6.8 33.3 matgl 0.0 4.5 14.6 matgl-direct 3.8 6.8 12.5 mattersim 0.0 4.5 16.7 orb-d3-v2 0.0 1.1 2.1 orb-v2 1.0 0.0 0.0 sevennet 0.0 2.3 41.7
@@
Table 2: The mean absolute error (MAE) and total computational time for a variety of material properties calculated with each uMLFF type. The MAE is calculated with respect to JARVIS-DFT data. Properties include (in order) lattice constants a, b, c, volume, bulk modulus, C11 and C44 components of the elastic tensor, and phonon band structure. The computational time is measured per CPU. An interactive and up-to-date version of the table is available in the JARVIS-Leaderboard.
@@
uMLFF Type err_a err_b err_c err_vol err_kv err_c11 err_c44 err_phonon Time Å Å Å Å3 GPa GPa GPa cm-1 seconds alignn_ff 0.087 0.102 0.137 9.95 122 196 74 150 8.8 x 104 chgnet 0.046 0.049 0.109 3.58 85 59 45 68 3.5 x 104 eqV2_153M_omat 0.036 0.040 0.111 2.90 117 70 32 49 1.6 x 102 eqV2_31M_omat 0.029 0.032 0.099 3.05 115 78 37 48 3.4 x 105 eqV2_31M_omat_mp_salex 0.028 0.030 0.096 3.17 112 47 41 48 2.7 x 105 eqV2_86M_omat 0.033 0.040 0.094 3.41 116 85 35 49 9.7 x 105 eqV2_86M_omat_mp_salex 0.027 0.030 0.100 3.14 114 45 35 48 7.2 x 105 mace 0.035 0.038 0.084 3.00 94 43 38 58 7.2 x 104 mace-alexandria 0.081 0.086 0.206 6.05 76 96 45 81 2.5 x 105 mace-mpa 0.044 0.046 0.098 3.57 110 35 34 49 8.4 x 104 matgl 0.052 0.057 0.128 3.50 83 64 41 78 1.7 x 104 matgl-direct 0.044 0.046 0.106 2.88 84 73 50 85 2.5 x 104 mattersim 0.031 0.033 0.110 3.18 110 32 32 48 5.9 x 104 orb-d3-v2 0.031 0.030 0.048 1.90 130 74 74 52 2.7 x 104 orb-v2 0.023 0.025 0.159 3.38 119 80 74 49 2.4 x 104 sevennet 0.035 0.038 0.100 3.43 103 46 37 53 5.7 x 104
@@
Figure 3: Parity plots for a) surface energy (in J/m2) and b) vacancy formation energy (in eV) for each uMLFF type with respect to JARVIS-DFT data (MAE is given in the inset). c) Depicts the parity plot for work of adhesion (Wad in J/m2) for several material interfaces and MAE with respect to previous calculated and experimental data (taken from Ref. 113). To the right of the parity plot, we focus on the best performing Wad prediction for the Al(111)Al2O3(001) interface, showing the xy grid search and optimal in-plane interface configurations with each uMLFF.
@@
Figure 4: The calculated radial distribution function ($g(r)$) as a function of pair separation distance computed with several uMLFF models for amorphous Si (by performing melt/quench simulations). Ab initio (AIMD) results are given as a ground truth benchmark (black dotted line) and the MAE and R2 with respect to AIMD is reported in the inset. The atomic structure for amorphous Si is given for AIMD and the best performing uMLFF models (MatterSim, eqV2_31M_omat and ORB).
@@
Fig. 1 depicts a full schematic of the CHIPS-FF workflow. In this workflow, initial structures are taken from the JARVIS-DFT database. JARVIS-Tools is used to pull structures from the JARVIS-DFT database and generate supercell surface and defect structures prior to atomistic simulations. Various uMLFF calculators are accessed through ASE. So far, we have implemented MatGL, CHGNet, ALIGNN-FF, MACE, SevenNet, ORB, MatterSim, and eqV2 (OMat24). As new pretrained force fields are developed over time, they will be added to the workflow. In addition, CHIPS-FF offers the flexibility for the user to add their own force field or property to the pipeline. Once each uMLFF is loaded and run through the ASE calculator, a variety of properties can be computed including: optimized structure, bulk modulus (equation of state fitting), elastic tensor (elastic package), phonons (phonopy), thermal expansion (phonopy), thermal conductivity (phono3py), molecular dynamics, vacancy formation energy, surface energy, and interface properties (InterMat). As these calculations are carried out, detailed logging information is saved, including the computational timing for each stage of the calculation and whether or not each stage converged. Once these material properties are computed with each uMLFF, they are cross checked with entries in the JARVIS-DFT database and the mean absolute error (MAE) for each property is computed. For spectral quantities such as the phonon band structure, we computed the MAE for each band along the q-point path using:
@@
For each of the 104 structures, we performed structural relaxations, fitting of the energy vs. volume curve (equation of state), calculation of the elastic tensor, calculation of the phonon band structure, relaxation of all nonpolar surfaces (among [1, 0, 0], [1, 1, 1], [1, 1, 0], [0, 1, 1], [0, 0, 1], and [0, 1, 0]) , and relaxation of each type of point defect in the material. The computational timing and errors in the lattice constants, elastic properties and phonons for each uMLFF model are depicted in Table 2. Out of the 104 materials in our test set, there were 85 matching entries for DFT phonon calculations in JARVIS, from which the composite MAE of the phonon band structure was computed in Table 2. For most structural and phonon properties, the OMat24 models perform exceedingly well, but are the most computationally intensive. Fitting the equation of state to obtain bulk modulus and calculating the elastic tensor prove to be a more challenging task for these uMLFF models, emphasizing difficulties in modeling the potential energy landscape far from equilibrium. We found that models such as MACE-Alexandria and MatGL have superior performance in predicting the energy vs. volume curves compared to the OMat and ORB models. For contributions to the elastic tensor (C11 and C44), we also observe relatively high errors. We see that only a few models can simultaneously predict C11 and C44 with reasonable accuracy (MAE less than 50 GPa). Interestingly, we find that eqV2_31M_omat, eqV2_86M_omat and eqV2_153M_omat can predict C44 with reasonable accuracy, but have much higher prediction errors for C11, while eqV2_31M_omat_mp_salex and eqV2_86M_omat_mp_salex do a better job at predicting C11 and C44 simultaneously. We observe that MACE-MPA-0 and MatterSim give the best results for simultaneous C11 and C44 predictions.
Excellent results for structural parameters and phonons are reported for orb-v2 and MatterSim at $\approx$ 10 - 100 times less the computational cost of the OMat24 models. ORB and MatterSim offer similar performance in terms of predicting phonon band structure when compared to several equivariant uMLFFs such as MACE, OMat24 and SevenNet. This is especially interesting since ORB and MatterSim-v1 are not equivariant uMLFFs. In addition, we see improved phonon results for the newer MACE-MPA-0 when compared to MACE-MP-0. Recently, it was reported that OMat and ORB models had substantial errors in phonon predictions when compared to models such as MatterSim, MACE, CHGNet, and SevenNet for 10,000 materials 90. It is entirely possible that ORB and OMat models can predict accurate phonons for our limited test set (104 materials) but fail for a broader and more diverse set of materials.
@@
In addition to summing the computational timing for the entire workflow, we also examined the scaling behavior of each uMLFF (on CPU). These results for supercells of Cu (up to 10,000 atoms) are depicted in Fig. 2a). Unsurprisingly, we see that OMat model scales much more drastically with system size as compared to the other uMLFF models. In addition, we find that the invariant models have more favorable scaling than the equivariant models such as MACE and SevenNet, with SevenNet having slightly better scaling than MACE. The scalability of ORB and MatterSim is a huge advantage due to their accuracy of various material property predictions. Similar scalability studies were conducted in Ref. 55 The quantities of surface energy and vacancy formation energy are difficult to predict with machine learning methods and provide a rigorous benchmarking test for uMLFF architectures. Out of the 104 materials, there exists 85 entries for surface DFT calculations and 48 entries for defect DFT calculations within JARVIS-DFT. We used these calculations to compare our uMLFF results. Fig. 3a) and Fig. 3b) depict the parity plots and corresponding MAE for surface energy and vacancy formation energy of our test sets. The very low error for surface energy (0.18 J/m2) and vacancy formation energy (0.94 eV) highlight some of the successes of recent state-of-the-art models such as OMat24, ORB, MACE-MPA-0 and MatterSim. We also observe generally good results for MACE-MP-0 and SevenNet when predicting vacancy formation energy and surface energies. The low cost of ORB, high rate of convergence and high accuracy make it a viable tool to relax larger surfaces and defect supercells.